OpenStudy (calculusxy):

Need help solving There are originally 1000 bacteria in a culture, and 4 hours later there are 4000. Find the rate of increase per hour of the bacteria. (Please do not use logs for this problem. I need the most basic way to do it with exponential growth.)

OpenStudy (calculusxy):

Should I use the formula: \(y = ab^x\)?

OpenStudy (jdoe0001):

sure \(\begin{array}{llrll} originally&1000&y = 1000(b)^o\\ 4hrs\ later&4000& y =1000(b)^x\\ &&4000=1000(b)^4\\ &&\cfrac{\cancel{4000}}{\cancel{1000}}=b^4\\ &&4=b^4\\ &&\sqrt[4]{4}=b \end{array}\)

OpenStudy (calculusxy):

That's what I did and got the answer of \(\approx\) 1.41. But the answer key says that the answer is 0.3467.

OpenStudy (jdoe0001):

\(\bf 1000( 0.3467)^4\approx 14.44\)

OpenStudy (jdoe0001):

not even close btw, so... can't be 0.3467

OpenStudy (jdoe0001):

now if you plug your 1.41, you'll get a workable number now, if you use 1.4142135623730950488 , you'll get pretty much 3999.99999999

OpenStudy (calculusxy):

That's what the answer key says. I tried it out with the formula of \(\large A = Pe^{r(t)}\) and it was 0.3467. But my teacher thinks that it would be too advanced to use that formula does not want me to use it. What can I do as an alternative?

OpenStudy (jdoe0001):

well, it doesn't fit for the equation of \(\bf y = 1000(b)^x\) though, 1.41 does

OpenStudy (jdoe0001):

or to be exact, 1.4142135623730950488 =)